The GPAs of entering freshmen at an elite university are known to be normally distributed with σ =.89. A random sample of 30 freshmen from the incoming class at a state university had a standard deviation of 1.7. What is the probability of getting a standard deviation of 1.7 or larger from a sample of size 3 if the distribution is normal with σ =.89?
Answer #1

( please revert if sample size is n=3 as in one place it is 30 and other it is 3 ; I am using 3 here)

as X2 test statistic =(n-1)s2./\sigma2

P(s>1.7)=P(S2 >2.89)=P((n-1)s2./\sigma2 >(3-1)*2.89/0.892)=P(X2 >7.297) =0.0260

(Note: